OpenStudy (anonymous):

y varies directly with x, and y = 2 when x = 5. What is the value of x when y = 15?

OpenStudy (shamil98):

y = kx 2 = 5k 2/5 = k 15 = (2/5)x solve for x.

OpenStudy (anonymous):

x=37.5 , thank you so much!

OpenStudy (anonymous):

y=kx. k=y/x. x=y/k k=2/5. x=15/0.4 k=0.4. x=37.5